OpenStudy (spring98):

(6^2)^x = 1, what is the value of x? Explain your answer!

OpenStudy (spring98):

@madhu.mukherjee.946

OpenStudy (spring98):

@Nnesha @sleepyjess @peachpi @mathlover2014

OpenStudy (spring98):

@misty_fox @

rishavraj (rishavraj):

hint a^0 = 1

rishavraj (rishavraj):

and (a^x)^y = a^{xy}

OpenStudy (spring98):

Can you explain this to me step by step i don't get this. plz

OpenStudy (spring98):

anyone?

OpenStudy (anonymous):

x = 0 because anything raised rto the 0 power is 1

OpenStudy (madhu.mukherjee.946):

6^2=36 so 36 ^0 =1

OpenStudy (spring98):

Can you also help me with this one plz? this is part B.

OpenStudy (madhu.mukherjee.946):

this is an indices property that anything raised to the power is 0

OpenStudy (spring98):

If (60)x = 1, what are the possible values of x? Explain your answer.

OpenStudy (madhu.mukherjee.946):

60x=1 x=1/60

OpenStudy (madhu.mukherjee.946):

this 60 multiplied to x and that was 36^x

OpenStudy (anonymous):

Assuming you mean 60^x=1, that's exactly the same question.

OpenStudy (madhu.mukherjee.946):

@Spring98

OpenStudy (madhu.mukherjee.946):

is it 60^x 0r 60 x

OpenStudy (spring98):

It is not the same question how do i do part B?

OpenStudy (madhu.mukherjee.946):

what is part b?

OpenStudy (madhu.mukherjee.946):

60x=1 x=1/60

OpenStudy (madhu.mukherjee.946):

@Spring98

OpenStudy (anonymous):

a. (6^2)^x = 1 b. 60^x = 1 These are the same problem with a different number in parentheses. If anything raised to the \(0\) power is 1, what does x have to be?

OpenStudy (spring98):

So is both Part A and Part B have the same answer?

OpenStudy (anonymous):

yes.

OpenStudy (madhu.mukherjee.946):

nope

OpenStudy (madhu.mukherjee.946):

part is (60)x=1 x=1/60

OpenStudy (spring98):

there can not be such answers in exponents or can there?

OpenStudy (anonymous):

@Spring98 yes there can be. that's called a rational exponent, my guess is you'll get there shortly. You should use 60^x to indicate exponents @madhu.mukherjee.946 was asking you to clarify the question above.

OpenStudy (anonymous):

60x = 1 is different from 60^x = 1

OpenStudy (madhu.mukherjee.946):

thats what i'm saying

OpenStudy (spring98):

sdo both the answers are 1 then right?

OpenStudy (spring98):

@peachpi @madhu.mukherjee.946

OpenStudy (anonymous):

The right side of all these equations are 1. The exponents are all 0. What are you looking for in your questions, the exponent or the right side?